SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (Oct. 25, 2005)

Attendees: Rainer Bartoldus, Prafulla Behera, Kris Chaisanguanthum, Al Eisner, Fang Fang, Selina Li, Rahmat Rahmat, Dong Su, Dan Walker.

Operations

  • No (cosmic) data taken yet as most detector components are still down.

EMT

  • Dan is getting ready and waiting for the EMC electronics to be back on to take some cosmics runs to make sure the new online release works with the EMT prescale in the configuration. The plan is to take some cosmics and then calibration runs. Note that the prescale should not be used in the calibration. Due to the EMC chiller problem today and computer outage at IR2 tomorrow, the ealiest we could get the EMT tests done is probably either Thursday or Friday of this week.
  • We would like to monitor the prescale bits, say, having a histogram with one bin for prescaled and other for un-prescaled events, etc.

DCT/DCZ

  • Another new version of the TSF firmware from Marc is being tested today. So far, it passed the autodetect (Dlink fiber removed/plugged in and master crate power cycled), MC algorithm and DAQ tests at the test stand. We are waiting for an opportunity to take some DCH-DCT calibrations when the DCH electronics is back to the test stand. If all go well, the next stage is to test it at IR2.

Online

  • Fang is working on improving the DCZ online code to save the highest bit corruption for debugging. The latest version of the code only skips the bad TSF segments rather than skipping the event as it was previously.
  • The new online release 13.7.0 based on offline 18.3.3 is now available and installed at IR2. Fang will make sure that the ROM lib will be updated before we take data again. This can wait till the next online release which should be available before data-taking resume in Nov.

GLT

  • Su Dong mentioned that for the single photon trigger, we could apply a L1 (line based) veto on 1E such as requiring #B<1 to try to get the rate go below 100Hz. Need to come up with an algorithm for L3 filter.

L3

  • Su Dong pointed out some problems with the L3 OPR plots that need to be fixed:
    • The x-axis for the track efficiency vs 1/Pt plot was cut off at +/-2 (should have been +/-5).
    • Why is the track efficiency vs phi only about 75%? Maybe the denominator has too loose requirements? It was suggested to check the tanlam or the Pt cut. Also check to see if they are independent of SVT tracks.
  • Rainer suggested maybe to fill these histograms with beta tracks.

Data Quality and Monitoring

  • Rahmat Rahmat from Oregon University is in training for Trigger QA.
  • The time history plots still have unresolved problems such as the first bin of the event rate is always way too high to reveal any structure with the rest of the bins. While we were looking at the time history plots in run 58712, we found that the number of TSF segments seem to be increasing with time within this run. This run is taken on Saturday Oct.8 when PEP was trying hard to reach 10^34 lumi. The runs before and after this run look normal for this plot though.
  • Su Dong mentioned that he will ask Brian to see if he can help fix some of these remaining problems with the time history plots. Rainer will also help look into the scalar problem.

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.